OpenStudy (anonymous):

35. Write an equation in point-slope form for the line through the given point with the given slope.(–7, 9); m = 4/5 (1 point) y + 7 = 4/5 (x – 9) y – 7 = 4/5 (x + 9) y – 9 = 4/5 (x + 7) y + 9 = 4/5 (x – 7)

OpenStudy (anonymous):

Your previous answer was correct I misplaced the answers though with another, it has been corrected.

OpenStudy (anonymous):

np, you use the point-slope form for the equation of a line: y - y1 = m(x - x1) -> y - 9 = (4/5)[x - (-7)] y - 9 = (4/5)(x + 7)

OpenStudy (anonymous):

Anything else I can do for you?

OpenStudy (anonymous):

lol I guess we're done then!

OpenStudy (anonymous):

I guess we are, have a nice day.

OpenStudy (anonymous):

u2, the best!